OpenStudy (anonymous):

g(x)= (5-x)/4 and h(x)= 8x-3 what is the simplest expression for gh(x)?

OpenStudy (anonymous):

g(8x-3)

OpenStudy (anonymous):

that's what you need to find

OpenStudy (anonymous):

g(8x - 3) = (5 -(8x - 3))/4

OpenStudy (anonymous):

= 8 - 8x/4

OpenStudy (anonymous):

= 2 - 2x

OpenStudy (anonymous):

thank you thats what i got as well was just checking